OpenStudy (anonymous):

Classify the solutions of 3 over x plus 5, plus one fifth, equals 2 over x plus 5 as extraneous or non-extraneous.

OpenStudy (anonymous):

x = -5; extraneous x = -5; non-extraneous x = -10; extraneous x = -10; non-extraneous ; x = -5; extraneous

OpenStudy (anonymous):

@Awesome781 @MegänD @SkiTTleoooo47 @cutegirl @minnie♥mouse @bowershe @Hannah28 @googlesun @goformit100 @jim_thompson5910

OpenStudy (anonymous):

@PurplePanda312 @Firebrand @Mr.ClayLordMath @asfasfasf @heather040200 @Mertsj @poopsiedoodle

OpenStudy (anonymous):

@Yuba @radar @Geniu @creamsupreme

OpenStudy (anonymous):

@Destinymasha @dee_faulkner @AskingAlexandria4life

OpenStudy (anonymous):

can somebody please help me

OpenStudy (anonymous):

let me see if i can read it

OpenStudy (anonymous):

\[\frac{3}{x+5}+\frac{1}{5}=\frac{2}{x+5}\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

this does have a solution, i can show you how to get it if you like

OpenStudy (anonymous):

yes that is and please do im so sorry for responding back late my little brother needed something @satellite73

OpenStudy (anonymous):

@radar

OpenStudy (radar):

 Well there is one root that is non-extraneous. lets look for more.